A simple random sample of 2500 BC single occupant dwellings revealed that 850 reported spending more than $750 a month on groceries.

a) Provide the point estimate of the population proportion of single occupant dwellings that spend more than $750 a month on groceries

b) Calculate the standard error of proportion using a minimum of 4 decimal places for intermediate calculations.
Answer #1

Here n = 2500 ; x = 850

(a) point estimate of the population proportion of single occupant dwellings that spend more than $750 a month on groceries = p^ = 850/2500 = 0.34

(b) Standard error = sqrt [p^ * (1-p^)/n] = sqrt [0.34 * 0.66/2500] = 0.0095